# FENDA-FL: Personalized Federated Learning on Heterogeneous Clinical Datasets

### v2

@@ Appendix E FENDA-FL Architecture Studies
Table 10: Results corresponding to different FENDA-FL architecture variations on the Fed-IXI task. “Depth (G, L)” and “Latent (G, L)” refer to the number of projection layers used in the global and local U-net feature extractors, respectively, along with the size of the latent space for each voxel.
fedixi_architecture_perturbations
@@
Table 10: Results corresponding to different FENDA-FL architecture variations on the Fed-IXI task. “Depth (G, L)” and “Latent (G, L)” refer to the number of projection layers used in the global and local U-net feature extractors, respectively, along with the size of the latent space for each voxel.
@@
Results for variations in the FENDA-FL model architecture for the Delirium task are reported in Table LABEL:delirium_architecture_perturbations. Aside from FENDA-FL, the best performing strategy for this task is FedProx with an accuracy of $0.8095$ and AUC/ROC of $0.8504$, using a model with $8,988,736$ parameters. From the table, it is evident that assigning more layers and a larger latent space to the global module is beneficial. Given that a global model is performing this task well, such an emphasis makes sense. However, the use of a, albeit smaller, personal module allows FENDA-FL to approach FedProx in terms of accuracy and surpass it in terms of AUC/ROC, despite having fewer parameters. Note that, in the imbalanced layer setting, the number of parameters lost, through reducing the depth of one module, are approximately transferred to the deeper module by increasing the layer sizes.
@@
Table LABEL:fedixi_architecture_perturbations shows results for FENDA-FL architecture perturbations for the Fed-IXI task. While the performance changes are fairly small, there is a slight advantage to using a larger local feature extraction module for this task. Doing so brings the FENDA-FL model results in line with the best performing strategy, APFL, for this problem. Note that a larger FENDA-FL model is used to allow for latent space imbalance in the U-Net. Note also that APFL does not appear to benefit from the additional parameters.
[…]
checkpoint_ablation_results Fed-Heart-Disease: Mean Accuracy Fed-IXI Optimal Parameters: Mean Dice Latest Global Local Latest Global Local FedAvg $0.706$ ($0.000$) $\mathbf{0.724}$ ($0.000$) $\mathbf{0.724}$ ($0.000$) $0.9845$ ($0.0002$) $0.9845$ ($0.0001$) $\mathbf{0.9846}$ ($0.0001$) FedAdam $0.711$ ($0.000$) $0.719$ ($0.000$) $\mathbf{0.742}$ ($0.000$) $\mathbf{0.9815}$ ($0.0001$) $0.9811$ ($0.0002$) $0.9811$ ($0.0002$) FedProx $0.703$ ($0.000$) $0.716$ ($0.000$) $\mathbf{0.721}$ ($0.000$) $0.9841$ ($0.0001$) $0.9842$ ($0.0001$) $\mathbf{0.9844}$ ($0.0001$) SCAFFOLD $\mathbf{0.717}$ ($0.000$) $0.711$ ($0.000$) $0.682$ ($0.000$) $\mathbf{0.9828}$ ($0.0000$) $0.9827$ ($0.0000$) $0.9827$ ($0.0000$) MOON $0.7029$ ($0.019$) $0.729$ ($0.009$) $\mathbf{0.741}$ ($0.009$) $0.9851$ ($0.0002$) $\mathbf{0.9852}$ ($0.0001$) $0.9851$ ($0.0002$) FedPer $\mathbf{0.814}$ ($0.003$) $\mathbf{0.814}$ ($0.002$) $\mathbf{0.9852}$ ($0.0003$) $0.9847$ ($0.0001$) APFL $0.777$ ($0.013$) $\mathbf{0.801}$ ($0.006$) $0.9858$ ($0.0000$) $\mathbf{0.9864}$ ($0.0002$) PerFCL $\mathbf{0.806}$ ($0.010$) $0.805$ ($0.018$) $0.9843$ ($0.0004$) $\mathbf{0.9847}$ ($0.0001$) FENDA-FL $0.802$ ($0.010$) $\mathbf{0.815}$ ($0.000$) $0.9848$ ($0.0026$) $\mathbf{0.9856}$ ($0.0001$) Fed-IXI Sub-Optimal Parameters: Mean Dice Latest Global Local FedAvg $0.9766$ ($0.0006$) $\mathbf{0.9788}$ ($0.0006$) $\mathbf{0.9788}$ ($0.0006$) FedAdam $0.9202$ ($0.0091$) $0.9483$ ($0.0056$) $\mathbf{0.9488}$ ($0.0054$) FedProx $0.9295$ ($0.0356$) $0.9548$ ($0.0131$) $\mathbf{0.9558}$ ($0.0136$) SCAFFOLD $\mathbf{0.4360}$ ($0.0000$) $\mathbf{0.4360}$ ($0.0000$) $\mathbf{0.4360}$ ($0.0000$) MOON $0.9177$ ($0.1155$) $0.9772$ ($0.0022$) $\mathbf{0.9777}$ ($0.0015$) FedPer $0.9754$ ($0.0019$) $\mathbf{0.9779}$ ($0.0015$) APFL $0.9780$ ($0.0007$) $\mathbf{0.9793}$ ($0.0005$) PerFCL $\mathbf{0.9853}$ ($0.0005$) $0.9711$ ($0.0282$) FENDA-FL $0.9705$ ($0.0075$) $\mathbf{0.9790}$ ($0.0004$)
Appendix F Federated Checkpointing Ablation Experiments
To quantify the benefits of the federated checkpointing strategies proposed in Section 3.3, we compare the results against the widely used fixed-rounds strategy, where only the model from the final FL round is saved. For non-personalized approaches, this is the final aggregated model on the server. For personalized approaches, the final-round client-side model is saved after any server-side aggregation. These checkpoints are referred to as “Latest” in Table LABEL:checkpoint_ablation_results.
For Fed-Heart-Disease, there is a clear advantage to federated checkpointing, especially the local version, for all strategies except of SCAFFOLD and PerFCL. Notably, APFL and FENDA-FL benefit significantly from local federated checkpointing, as do most of the non-personalized methods. For Fed-IXI, two experiments are conducted. In the first, the optimal hyper-parameters for each strategy are applied during training. The benefits of local checkpointing are less pronounced in this setting, but some of the approaches exhibit small improvements. APFL, PerFCL, and FENDA see larger improvements. In examining the training trajectories, many of the aggregated validation losses do not show signs of overfitting, likely reducing the advantage of federated checkpointing.
To consider settings when the optimal number of rounds is not well calibrated, training with sub-optimal hyper-parameters for each strategy is performed. For all strategies, except PerFCL, a client-side learning rate of $0.1$ is applied. For FedAdam and SCAFFOLD, a server-side learning rate of $0.1$ and $1.0$, respectively, is used. FedProx sets $\mu$ = $1\mathrm{e}{\text{-}3}$, APFL applies a learning rate for $\alpha$ of 1.0, and MOON sets $\mu$ to $1\mathrm{e}{\text{-}3}$. Finally, PerFCL uses a client-side learning rate of $1\mathrm{e}{\text{-}3}$, $\mu=0.01$, and $\gamma=1$. In this context, a significant advantage is seen when using federated checkpointing, with the exception of PerFCL. Local federated checkpointing also outperforms the server-side analogue. Observe also that federated checkpointing results in a measurable reduction in variance for all strategies, again excepting PerFCL. The large performance drop and variance for PerFCL with federated checkpointing can be traced to a single federated run failing to converge during training, yielding poor performance.
[…]
Figure 4: Changes in the average metric attained by models trained with APFL when the learning rate for $\alpha$ is varied. The method is fairly robust to such changes, but there is still a measurable impact on the performance when such learning rates are poorly chosen.
Appendix G APFL LR Studies
The stability of the APFL algorithm with respect to the choice of the learning rate for $\alpha$ across each of the FLamby datasets is considered in this section. Figure LABEL:apfl_alpha_lr_studies shows the average metric across five APFL training runs when the learning rate for $\alpha$ is varied. In all cases, $\alpha$ is initially set to $0.5$. The performance of models trained with APFL is fairly robust to variations in the learning rate across each dataset, especially Fed-IXI. However, an impact on accuracy and balanced accuracy for the Fed-Heart-Disease and Fed-ISIC2019 tasks is observed. In both settings, a full percentage point is lost for poorly chosen learning rates compared to the optimal learning rate.
@@ Appendix H Momentum-Based Server-Side Optimizers and BN
Certain layers in deep learning models accumulate state throughout training that must remain strictly positive. Batch normalization (BN) (Ioffe and Szegedy, 2015) is one of these layers. During training, the layer accumulates mean and variance estimates of preceding layer activations. These estimates are used to perform normalization during inference. When using FedAvg, these estimates are aggregated, like all other parameters, using linearly-weighted averaging. Thus, strictly positive local quantities result in strictly positive aggregated quantities. However, when using momentum-based server-side optimization, as in FedAdam, this property is no longer preserved. This is particularly problematic for BN, where the square root of the variance is computed to perform normalization.
@@
As an illustration of this phenomenon, consider learning a single parameter $x$ and an initial guess of $x_{0}=2.0$. Set the FedAdam parameters to $\beta_{1}=0.9$, $\beta_{2}=0.9$, $\tau=1\mathrm{e}{\text{-}9}$, and a learning rate of $\eta=0.1$. Assume that each client calculates a new parameter of $\tilde{x}=0.1$ at each iteration. Algorithm 1 shows the iterative calculations of FedAdam for this simplified setting. Running these iterations results in a negative value for $x_{t}=-0.204$, despite $\tilde{x}>0$ for every iteration. This drift is due to the momentum associated with the updates. Therefore, in settings where BN is present and the FedAdam strategy is employed, batch tracking is shut off to avoid such issues. It should be noted that this is qualitatively similar to the FedBN strategy in (Li et al., 2021b).
[…]
delirium_extreme_heterogeneity Refer to caption
Appendix I Extreme Heterogeneity
As a test of the robustness of the FL methods studied here, with particular interest in FENDA-FL, an experiment inducing extreme feature-space heterogeneity is designed. For the GEMINI Delirium task, described in Section 3.2, each client further processes its data with dimensionality reduction through PCA and feature selection to produce input vectors of dimension $300$. As the clients independently perform this process, the resulting feature spaces are no longer well-aligned. In this setup, locally trained models are unlikely to generalize well to other domains. Moreover, FL techniques that produce a single shared model are expected to encounter convergence issues. This is confirmed in the results shown in Figure LABEL:delirium_extreme_heterogeneity. Each client’s local model generalizes poorly to other domains and global FL approaches struggle with performance. On the other hand, APFL and FENDA-FL fare better, as they are able to lean on local training when global parameters are not informative. FENDA-FL demonstrates notable robustness to this setting, recovering the siloed performance target. Note also that local checkpointing is important for this task.
Appendix J FENDA-FL Details and Distribution Shift
Assume there is a central server and $M$ clients, each with a local dataset $D_{i}$, for $i\in\{1,\dots,M\}$. Client data exists on domain $X\times Y$ where $X\in\mathbb{R}^{I}$ and $Y\in\mathbb{R}^{O}$. FENDA-FL uses the FedAvg algorithm as a foundation to enable collaborative learning among clients. In particular, along with a global feature extractor $f_{{\mathbf{w}}}:X\rightarrow\mathbb{R}^{H_{G}}$, each client, $i$, has a local feature extractor $f_{{\mathbf{w}}_{i}}:X\rightarrow\mathbb{R}^{H_{i}}$ and classification head $g_{i}:\mathbb{R}^{(H_{G}+H_{i})}\rightarrow Y$. For a given input sample ${\mathbf{x}}\sim P_{i}(X)$, $f$ and $f_{i}$ map ${\mathbf{x}}$ to representations, ${\mathbf{z}}$ and ${\mathbf{z}}_{i}$, that are concatenated and passed into $g_{i}$ to yield a prediction, $\hat{y}_{i}$. In the simplest case, $g_{i}$ is a fully connected layer parameterized by weights $\theta_{i}\in\mathbb{R}^{O\times(H_{G}+H_{i})}$ and an activation $\sigma$ such that
$\hat{y}_{i}=\sigma\left(\theta_{i}\begin{bmatrix}{\mathbf{z}}\\ {\mathbf{z}}_{i}\\ \end{bmatrix}\right)=\sigma\left(\theta_{i}\begin{bmatrix}f_{{\mathbf{w}}}({\mathbf{x}})\\ f_{{\mathbf{w}}_{i}}({\mathbf{x}})\\ \end{bmatrix}\right).$
The global feature extractor weights, ${\mathbf{w}}$, are aggregated across clients via FedAvg, whereas the local feature extractor and classification head weights, ${\mathbf{w}}_{i}$ and $\theta_{i}$, are exclusively learned through SGD, or its variants, on each client. Both global and local parameters are iteratively updated across FL rounds.
J.1 Distribution Shifts in FL
@@ Appendix I Extreme Heterogeneity
Distribution shifts arise when the joint distribution of inputs and targets varies between datasets (Quinonero-Candela et al., 2008). In the standard ML setup, distribution shifts occur between the train and test data, such that $P_{\text{train}}(X,Y)\neq P_{\text{test}}(X,Y)$, leading to poor model generalization. The most common distribution shifts include label, covariate and concept drift (Zhang et al., 2021). Label shift is characterized by different label distributions $P(Y)$, but constant class conditional distribution, $P(X|Y)$. Covariate shift occurs when the marginal distribution, $P(X)$, varies across datasets while the conditional distribution, $P(Y|X)$, remains the same. Finally, concept drift arises when the class conditional distribution $P(Y|X)$ varies and $P(Y)$ remains fixed.
@@
In the FL setting, distribution shifts are common among clients, due to variations in the underlying data generating process. These variations may stem, for example, from differences between the sensors, preferences, geographies and time periods involved in data collection (Kairouz et al., 2021). In this regime, FedAvg often exhibits slow convergence and poor performance. Many works have attempted to address this for various types of shifts (Karimireddy et al., 2020; Deng et al., 2020; Li et al., 2021b).
@@
As a personalized method, FENDA-FL allows each client to train a custom model while maintaining the benefit of using FL to jointly learn a subset of weights. In particular, as described above, FENDA-FL has a locally-optimized feature extractor and classification head, in addition to a global feature extractor learned via FedAvg. The local feature extractor and classification head endow client models with a degree of resilience to label and covariate shift. They may adapt feature representations and predictions to lean more or less heavily on global representations or modify prediction distributions to match local statistics. Beyond such shifts, FENDA-FL is able to handle concept drift where the conditional distribution, $P(Y|X)$, varies across clients. For example, in the extreme case, this is accomplished by the local classification head ignoring global features altogether. This is not true of non-personalized FL models, which assume a single function exists that obtains low error across all clients.
@@
FENDA-FL is also resilient to extreme distribution shifts involving misalignment of client features. As demonstrated in Appendix I, FENDA-FL is able to obtain performance consistent with locally trained models when the feature spaces of the clients are not aligned, where as other benchmark approaches fall significantly short. One type of distribution shift that is not admissible in the proposed approach is misaligned label spaces. This is an important setting to explore in future work.

### v3

@@ Appendix F FENDA-FL Architecture Studies
Table 12: Results corresponding to different FENDA-FL architecture variations on the Fed-IXI task. “Depth (G, L)” and “Latent (G, L)” refer to the number of projection layers used in the global (G) and local (L) U-Net feature extractors, respectively, along with the size of the latent space for each voxel.
@@
Figure 4: Generalization measurements for a selection of FL and non-FL approaches. Rows correspond to a model trained using either client-only data, centralized data, or with the indicated FL method. Columns denote the average performance of the model on the local test data of the named client.
@@
Results for variations in the FENDA-FL model architecture for the Delirium task are reported in Table 11. It is evident that assigning more layers and a larger latent space to the global module is beneficial. Given that FedProx, a global approach, is performing this task well, see Table 4, such an emphasis makes sense. However, the use of an, albeit smaller, personal module allows FENDA-FL to surpass FedProx in terms of AUC/ROC and be competitive with respect to accuracy, despite having fewer parameters.
@@
Table 12 shows results for FENDA-FL architecture perturbations for the Fed-IXI task. While the performance changes are modest, there is an advantage to using a larger local feature extraction module for this dataset. Doing so brings the FENDA-FL model results ahead of the previously best performing strategy, APFL, for this problem. Note that a larger FENDA-FL model is used to allow for latent space imbalance in the U-Net, but that APFL does not benefit from such an increase.
Appendix G Continued Analysis of Fed-ISIC2019 Results
For Fed-ISIC2019, many of the personalized FL approaches, excluding Ditto, produced lower than expected performance, see Table 2. Some insights can be found in Figure 4. SCAFFOLD, the best non-personalized FL method, outperforms the personalized models on test data drawn from Client 1 and Client 2. This is reversed when considering Client 4. However, the most significant difference is that most personalized models perform quite poorly on the test data of Client 5. Strikingly, Ditto does not suffer from the same precipitous drop, performing well across all client data.
Notably, Client 5 has the smallest dataset by a large margin; see Table 8. Moreover, the data appears to be of poor quality, with locally trained models only achieving a balanced accuracy of $0.425$. Alternatively, models trained only on data from Client 0 generalize well to Client 5. We hypothesize that the low quality of Client 5’s data impedes FL model training for most personalized methods. Further, we posit that SCAFFOLD’s control-variate corrections and the strong regularization of Ditto to the global model provide resilience to such local client drift. In future work, we aim to consider ways of overcoming this issue for other methods. Some potential avenues include additional loss terms to ensure the global and local feature extractors remain individually useful for classification or using a global FL approach to initialize model components as a warm start.
@@ Appendix H Momentum-Based Server-Side Optimizers and BN
Certain layers in deep learning models accumulate state throughout training that must remain strictly positive. Batch normalization (BN) (Ioffe and Szegedy, 2015) is one such layer. During training, the layer accumulates mean and variance estimates of preceding layer activations. These estimates are used to perform normalization during inference. When using FedAvg, these estimates are aggregated, like all other parameters, using weighted averaging. Thus, strictly positive local quantities result in strictly positive aggregated quantities. However, when using momentum-based server-side optimization, as in FedAdam, this property is no longer preserved. This is particularly problematic for BN, where the square root of the variance is computed to perform normalization.
@@
As an illustration of this phenomenon, consider learning a single parameter $x$ from an initial guess of $x_{0}=2.0$. Set the FedAdam parameters to $\beta_{1}=0.9$, $\beta_{2}=0.9$, $\tau=1\mathrm{e}{\text{-}9}$, and a learning rate of $\eta=0.1$. Assume that each client calculates a new parameter of $\tilde{x}=0.1$ at each iteration. Algorithm 1 shows the iterations for this simplified setting. After $30$ steps, $x_{t}=-0.204$, despite $\tilde{x}>0$ for every iteration. This drift is due to the momentum associated with the updates. Therefore, in settings where BN is present and the FedAdam strategy is employed, batch tracking is shut off to avoid such issues. It should be noted that this is qualitatively similar to the FedBN strategy in (Li et al., 2021c).
Appendix I FENDA-FL: Distribution Shift and Domain Adaptation
I.1 Distribution Shifts in FL
@@
Distribution shifts arise when the joint distribution of inputs and targets varies between datasets (Quinonero-Candela et al., 2008). For standard ML, shifts occur between the train and test data, such that $P_{\text{train}}(X,Y)\neq P_{\text{test}}(X,Y)$, leading to poor generalization. The most common shifts include label, covariate, and concept drift (Zhang et al., 2021). Label shift is characterized by different label distributions, $P(Y)$, but constant class conditioned distributions, $P(X|Y)$. Covariate shift occurs when the marginal distribution, $P(X)$, varies across datasets while the conditional distribution, $P(Y|X)$, remains the same. Concept drift arises when the class conditional distribution $P(Y|X)$ varies and $P(Y)$ remains fixed.
@@
In FL, distribution shifts are common among clients, due to variations in the underlying data generating process. These variations may stem, for example, from differences between sensors, geographies and time periods involved in data collection (Kairouz et al., 2021). In this regime, FedAvg often exhibits slow convergence and poor performance. Many works have attempted to address this for various specific types of shifts (Karimireddy et al., 2020; Deng et al., 2020; Li et al., 2023).
@@
As a personalized method, FENDA-FL allows each client to train a custom model while maintaining the benefit of using FL to jointly learn a subset of weights. In particular, FENDA-FL has a locally-optimized feature extractor and classification head, in addition to a global feature extractor learned via FedAvg. The local feature extractor and classification head endow client models with a degree of resilience to label and covariate shift. They may adapt feature representations and predictions to lean more or less heavily on global representations or modify prediction distributions to match local statistics. Beyond such shifts, FENDA-FL is able to handle concept drift where the conditional distribution, $P(Y|X)$, varies across clients. For example, in the extreme case, this is accomplished by the local classification head ignoring global features altogether. This is not true of non-personalized FL models, which assume a single function exists that obtains low error across all clients.
@@
FENDA-FL is also resilient to extreme distribution shifts. As demonstrated in Section 5.3, FENDA-FL is able to obtain improved performance compared to locally trained models even when the feature spaces of the clients are not aligned, whereas other benchmark approaches fall short. However, one type of distribution shift that is not admissible in the proposed approach is misaligned label spaces. This is an important setting to explore in future work.
I.2 Domain Adaptation and Generalization
FENDA-FL is connected to Domain Adaptation (DA). Generally, DA involves learning a model from source datasets that generalizes to target datasets. During training, subsets of the target datasets, typically unlabeled or sparsely labeled, are available. Federated DA considers the setting of distributed source and target datasets. A server may host the source dataset coupled with client-based, unlabelled datasets (Yao et al., 2022) or vice-versa (Peng et al., 2019). Alternatively, both source and target datasets may be distributed (Shen et al., 2023b; Song et al., 2020). Similarly, Federated Domain Generalization focuses on the task of training a model on distributed, private source domains that generalizes to unseen target domains (Li et al., 2023). Such methods focus on training models for out-of-distribution clients rather than (personalized) models optimized for known client data. These settings are related to those considered here but have different objectives and data constraints. However, FENDA-FL establishes a novel and useful conceptual bridge between FL and DA.